Q 1 Below are some data from the land of milk and honey. 

Year   Price of Milk  Quantity  of Milk   Price  of Honey  Quantity  of Honey 

2016   $1            100 quarts         $2                50 quarts 

2017   1              200              2               100 

2018   2              200              4               100


Compute nominal GDP, real GDP, and the GDP deflator for each year, using 2016 as the base year.




1
Expert's answer
2021-03-16T11:15:08-0400

Let's calculate the real GDP in 2016 year:


"Real\ GDP_{2016}=100\ quarts\cdot\dfrac{\$1}{quart}+50\ quarts\cdot\dfrac{\$2}{quart}=\$200."

Since 2016 is the base year, the real GDP equals the nominal GDP:


"Real\ GDP_{2016}=Nominal\ GDP_{2016}=\$200."

Let's calculate the real GDP in 2017 year:


"Real\ GDP_{2017}=200\ quarts\cdot\dfrac{\$1}{quart}+100\ quarts\cdot\dfrac{\$2}{quart}=\$400."

Let's calculate the nominal GDP in 2017 year:


"Nominal\ GDP_{2017}=200\ quarts\cdot\dfrac{\$1}{quart}+100\ quarts\cdot\dfrac{\$2}{quart}=\$400."


Let's calculate the real GDP in 2018 year:


"Real\ GDP_{2018}=200\ quarts\cdot\dfrac{\$1}{quart}+100\ quarts\cdot\dfrac{\$2}{quart}=\$400."

Let's calculate the nominal GDP in 2018 year:


"Nominal\ GDP_{2018}=200\ quarts\cdot\dfrac{\$2}{quart}+100\ quarts\cdot\dfrac{\$4}{quart}=\$800."

The GDP deflator can be calculated as follows:


"GDP\ deflator=\dfrac{Nominal\ GDP}{Real\ GDP}\times100."

Let's calculate the GDP deflator for each year:


"GDP\ deflator_{2016}=\dfrac{\$200}{\$200}\times100=100,""GDP\ deflator_{2017}=\dfrac{\$400}{\$400}\times100=100,""GDP\ deflator_{2018}=\dfrac{\$800}{\$400}\times100=200."